an airplane flew for 4 hours against a headwind of 40km/h. on its return the same wind was a tailwind and the flight took 3 hours. whats the planes speed in still air?? what two system of equations could i do?
Answer by stanbon(75887)

an airplane flew for 4 hours against a headwind of 40km/h. on its return the same wind was a tailwind and the flight took 3 hours. whats the planes speed in still air?? what two system of equations could i do?
-----------
Let s be the speed of the plane in still air and d be the distance flown.
WITH the wind DATA:
Distance = d km ; time=3 hrs ; rate= p+40 km/h
----------
AGAINST the wind DATA:
Distance = d km ; time=4 hrs ; rate = p-40 km/hr
--------
EQUATION:
distance = rate*time
1st: d=(p+40)3
2nd: d=(p-40)4
Then:
(p+40)3=(p-40)4
3p+120=4p-160
p=280 km/hr
--------
Cheers,
Stan H.